OpenStudy (anonymous):

What is a zero of the function f(x)= 2x-10

OpenStudy (anonymous):

to find the zero of a polynomial u have to find x such that f(x)=0

OpenStudy (anonymous):

so put f(x) =0 and solve for x.that will give you zero of f(x)

OpenStudy (anonymous):

so it will be 10 ?

OpenStudy (anonymous):

no u put f(x)=0 this is the same as 2x-10=0 because f(x)=2x-10 now take 10 to the right hand side u get 2x=10 now think what would x be?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

yes you are right

OpenStudy (anonymous):

I have another question

OpenStudy (anonymous):

so x=5 is a zero of f(x)

OpenStudy (anonymous):

please feel free to ask

OpenStudy (anonymous):

What is the solution to the system of equations above ? x - y = 5 x + y = 7

OpenStudy (precal):

you are looking for the point of intersection in this case

OpenStudy (anonymous):

so could you use the elimination method ?

OpenStudy (anonymous):

yes use elimination

OpenStudy (precal):

actually just add both equations together (addition method) 2x=12 so what is x?

OpenStudy (anonymous):

I was thinking about that , and x would be 6 .

OpenStudy (precal):

yes not sub x=6 into one of the two equations (either one will work) what is y?

OpenStudy (anonymous):

1 ! (:

OpenStudy (precal):

yes

OpenStudy (anonymous):

I have more questions , can I ask them ?

OpenStudy (precal):

post them separately please. Close this question and post a new one. If I have time I will help you, I have to do my homework

OpenStudy (anonymous):

ok